OpenStudy (tashia):

the vertical distance from the bottom of a staircase to the top is 20 feet. the angle of elevation is 25. find the distance from the first step to the last (hypotenuse).

OpenStudy (anonymous):

use trignometric ratios

OpenStudy (anonymous):

cos(25)=20/x ---> x=20/cos(25)